OpenStudy (anonymous):

How do I determine quadratic function f, whose graph is vertex (1,-2) and passes thru (3,6)

OpenStudy (anonymous):

Let the quadratic function be of the genera form \[ f(x)=ax^2+bx+c\] Then match the three unknown coefficients to the points \[f(1)=-2 \\ f(3)=6\] And for the Vertex \[1 = \frac{-b}{2a} \rightarrow 2a=-b \] Then solve system of equation

OpenStudy (anonymous):

ok, but how do I plug my numbers in

OpenStudy (anonymous):

\[f(1)=a+b+c=-2 \\ f(3)=9a+3b+c = 6 \\2a=-b \]

OpenStudy (anonymous):

System of equations, three unknowns, three equations, therefore one definite solution, because none of the equations above is redundant.

OpenStudy (anonymous):

i still don't understand

OpenStudy (anonymous):

Do you know how to solve a system of linear equations?